OpenStudy (anonymous):

Simplify sqrt(20) x 4sqrt(45)

OpenStudy (anonymous):

\[\huge \sqrt {20} \times 4 \sqrt{45} =\sqrt {4 \times 5} \times 4 \sqrt{9 \times 5} \] \[\huge =\sqrt {2^2 \times 5} \times 4 \sqrt{3^2 \times 5} \] \[\huge =2\sqrt {5} \times 4 \times 3 \sqrt{5} \] \[\huge =2\sqrt {5} \times 12 \sqrt{5} = 2 \times 12 \times \sqrt {5 \times 5}\] \[\huge = 24\times \sqrt {25} = 24 \times 5 = 120\]